OpenStudy (anonymous):

which eponentioal funtion has a y- intercept of 9, a horizonal asmtote of y=5 and contains the point (1,17) ? a f(x)-4*3^x+5 B f(x)=3*4^x+5 C f(x)=5*4^x+3 D f(x)=3*5^x+4

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

y intercept = 9 so if you plug in x=0, you should get y = 9 in your equation. for which option is this true?

OpenStudy (anonymous):

???

hartnn (hartnn):

consider last option, 3*5^x +4 if I put x= 0, i get 3*5^0 +4 = 3*1 + 4 = 7 and not 9. so last option cannot be the answer.

hartnn (hartnn):

can you similarly check other options?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait yea c ??

OpenStudy (anonymous):

is it c

hartnn (hartnn):

nope, 5+3= 8

hartnn (hartnn):

neither B 3+5 = 8

hartnn (hartnn):

so only A remains :)